# Approximate Vector Set Search: A Bio-Inspired Approach for High-Dimensional Spaces

### v1

@@ Abstract
Vector set search, an underexplored similarity search paradigm, aims to find vector sets similar to a query set. This search paradigm leverages the inherent structural alignment between sets and real-world entities to model more fine-grained and consistent relationships for diverse applications. This task, however, faces more severe efficiency challenges than traditional single-vector search due to the combinatorial explosion of pairings in set-to-set comparisons. In this work, we aim to address the efficiency challenges posed by the combinatorial explosion in vector set search, as well as the curse of dimensionality inherited from single-vector search. To tackle these challenges, we present an efficient algorithm for vector set search, BioVSS (Bio-inspired Vector Set Search). BioVSS simulates the fly olfactory circuit to quantize vectors into sparse binary codes and then designs an index based on the set membership property of the Bloom filter. The quantization and indexing strategy enables BioVSS to efficiently perform vector set search by pruning the search space. Experimental results demonstrate over 50 times speedup compared to linear scanning on million-scale datasets while maintaining a high recall rate of up to 98.9%, making it an efficient solution for vector set search.
@@
I Introduction
Vector search is a fundamental computational problem in various domains such as information retrieval [18], recommender systems [31], and computer vision [21]. The majority of vector search methods, tailored for querying by a single vector, address the curse of dimensionality [19] and the sparsity issue in high-dimensional vector space [50]. In this paper, we study a novel and challenging problem called Vector Set Search. Given a query vector set $\mathbf{Q}$, the objective is to find the top-$k$ most relevant vector sets from a vector set database $\mathbf{D}=\{\mathbf{V}_{1},\mathbf{V}_{2},\dots,\mathbf{V}_{n}\}$. This problem transcends the boundaries of single-vector search [12] by incorporating matching and aggregation between vectors within each set. This introduces additional computational complexity, for which there are currently no effective solutions.
Applications. Vector set search has many applications due to its structural alignment with real-world entities, for example:
  • Academic Entity Discovery. Academic entity discovery facilitates effective exploration of the rapidly growing literature. By representing academic entities as vector sets, such as a paper’s profile [48] consisting of vectors representing its citations, or an author’s profile [38] as a set of vectors representing their publications, vector set search enables effective identification of relevant research and tracking of academic trends.
  • Multi-Modal Query Answering. Multi-modal query answering systems leverage vector set search to enhance accuracy. As demonstrated in [43], this approach encodes multi-modal data (e.g., images, text, and audio) into vector sets using a multi-modal encoder [26]. Vector set search then enables accurate queries across multi-modal representations, improving the overall performance of multi-modal query answering systems.
  • Multi-Vector Ranking for Recommender. Multi-vector ranking enhances recommendation systems by representing users with vector sets. This approach, as demonstrated in [30], represents users as vector sets to capture various aspects of user profiles, including browsing history, search queries, click records, etc. The ranking process compares these vector sets, enabling flexible matching between diverse user interests.
@@ I Introduction
Let us delve into scholar search, an exemplar of vector set search that serves as a key focus in our experimental study.
Example 1: Figure 1 illustrates an exemplar data flow of vector set search applied to scholar search [38], preceded by vector set database preparation. The vector set database construction involves representing scholar profiles as vector sets derived from their publications [9], creating the database of scholars. The vector set search then operates on this database, comprising two key stages: distance computation and neighbor identification. The distance computation stage involves calculating the pairwise distances (such as Euclidean distance) between the query vector set and another set in the database, followed by aggregating (such as combining $max$ and $min$) into set-to-set distances. The neighbor identification phase then determines the top-$k$ nearest scholars based on these set distances for the query. This process enables the discovery of similar scholars for subsequent research activities.
Motivation. The example above reveals a key insight: the structure of vector sets aligns well with real-world entities, enabling effective applications. Recent advancements in embedding models [9] that generate high-dimensional, semantically rich vectors have further expanded the applicability. However, the computational demands of high-dimensional vectors and the combinatorial explosion of pairwise comparisons [3] complicate the development of efficient search.
Recent relevant studies such as DESSERT [11] utilize hash tables to accelerate specific non-metric distance computation. However, the fundamental requirements for vector set similarity assessment — precise differentiation capability and consistent measurement across comparisons — demand careful consideration of distance metrics (detailed analysis in Appendix VII-A). The Hausdorff [16, 17] distance naturally aligns with these requirements, providing mathematically rigorous set-to-set similarity measurement while maintaining essential symmetric properties for reliable similarity assessment.
Motivated by these fundamental requirements for robust set comparison, we recognize that metric distances are crucial for consistent evaluations. The Hausdorff distance, a well-studied metric in set theory [16, 17], offers native support for set distance measurement. However, it requires complex pairwise distance calculations, making it computationally intensive. This computational burden underscores the urgent need for more efficient vector set search methods.
Challenges. To provide efficient vector set search in high-dimensional spaces using Hausdorff distance, several key challenges arise. The first challenge is the curse of dimensionality [19]. Recent methods address this challenge through hash table construction [11] and the transformation of vector set to single vector [23]. However, both methods are limited to specific non-metric distances, constraining their applicability in Hausdorff distance. The second challenge is the Aggregation complexity. Existing Hausdorff distance search algorithms [44, 1, 32] leverage geometric information to decrease the number of aggregation operations. However, geometric information loses effectiveness in high dimensions.
Contributions. To overcome the above two challenges, we propose an algorithm to support an approximate top-$k$ vector set search. Specifically, our work draws inspiration from a locality-sensitive hashing [8, 37] inspired by the olfactory circuit [27]. We simulate the olfactory circuit to quantize vectors into sparse binary codes, addressing the curse of dimensionality. Based on this, an index is designed utilizing the set membership property of Bloom filter, which reduces the search space and decreases the number of aggregation operations. Overall, our contributions are summarized as follows:
@@
  • We define the first approximate vector set search problem in high-dimensional spaces using Hausdorff distance, which is a native set metric distance (see Section III).
@@
  • We propose BioVSS, which uses locality-sensitive property to accelerate vector set search, and provide a comprehensive theoretical analysis with proofs validating the correctness of the proposed method (see Section IV).
@@
  • We present an enhanced version of BioVSS called BioVSS++, which employs a dual-layer cascaded filter composed of inverted index and vector set sketches to reduce unnecessary scans (see Section V).
@@
  • We conduct extensive experiments showing our method achieves over 50 times speedup compared to linear scanning on million-scale datasets while maintaining a recall of up to 98.9%, validating its efficiency (see Section VI).
II RELATED WORK
@@
Single-Vector Search.  Single-vector search [49] is increasingly becoming a unified paradigm with the development of embedding models. However, efficiency remains a critical challenge for practical application. Various approximate search algorithms [25] have been developed to perform efficient single-vector search. These algorithms fall into three main categories: locality-sensitive hashing methods [37, 37, 47, 51], graph-based methods [28], and space partition methods [24, 15]. Locality-sensitive hashing like FlyHash [37] maps similar vectors to the same hash buckets. Graph-based approaches such as HNSW [28] construct navigable graphs for efficient search. Space partitioning methods, such as IVFFLAT [49, 10], IndexIVFPQ [14, 20, 10] and IVFScalarQuantizer [10], utilize k-means [46] clustering to partition the vector space and build inverted file indices for fast search. However, existing approaches focus solely on single-vector rather than set-based queries.
@@
Vector Set Search.  The problem of vector set search remains relatively unexplored and has not received sufficient attention in the research community. Early research focused on low-dimensional vector space, treating vector set search as a set matching problem often solved with the Kuhn-Munkres algorithm [22]. In geospatial research, trajectories were typically represented as point collections and measured with the Hausdorff distance [2, 44, 45]. Researchers developed various optimizations, such as incorporating road network information [36], using branch-and-bound techniques [32], and employing estimated Hausdorff distances for accelerated computations [1]. However, these approaches, tailored for low-dimensional spaces, do not extend well to high-dimensional spaces. Recent advancements in embedding models enable high-dimensional vectors to carry rich semantic information. By transforming vector sets into single vectors [23], existing high-dimensional single-vector approximate search algorithms can be leveraged to speed up the computational process. However, this transformation relies on a specific distance metric proposed by the work. While the method proposed in [11] is an efficient algorithm that accelerates the vector set search process through hash table construction, it has limitations in terms of measures. Specifically, this approach lacks theoretical support for the Hausdorff distance metric, particularly when the $max$ function is used as the outer aggregation strategy. Moreover, these methods lack effective support for Hausdorff distance. To address this gap, we focus on high-dimensional vector set search using Hausdorff distance.
Bio-Inspired Hashing & Bloom Filter.  Olfactory systems in various species show remarkable precision in identifying odorants [6]. For instance, Mice (rodent animals) analyze neural responses in the olfactory bulb to discriminate odors [35]. Similarly, Zebrafish (teleost animals) detect water-dissolved amino acids for olfactory signaling [33]. Furthermore, the olfaction of fly (insect animals) research is the most in-depth and has led to the development of a sparse locality-sensitive hashing function [8, 37] through computational simulations. Fly-inspired hashing [8], a novel locality-sensitive hashing function for single vector search, draws inspiration from the fly olfactory circuit. In the olfactory circuit, only a few neurons respond to specific odor molecules, creating sparse activation patterns that enhance odor recognition efficiency. BioHash [37] improves accuracy by learning the intrinsic patterns of data to set the connection strengths of projection neurons in the fly olfactory circuit [42]. Similarly, a Bloom filter maps set elements into a sparse array using hash functions, enabling efficient set storage. The binary Bloom filter [42] includes a binary array, where each position is either 0 or 1. Additionally, the count Bloom filter [5] offers finer granularity through counting. This work leverages the locality-sensitive property of the fly olfactory circuit and the set-storing capabilities of Bloom filters. Meanwhile, we exploit their structural similarities to construct an efficient index.
III DEFINITIONS & PRELIMINARIES
@@
To elucidate the computational process of Hausdorff distance, let us consider a concrete example.
Example 2: The computation process of Hausdorff distance is illustrated in Figure 2. Given two finite vector sets $\mathbf{Q}=\{\mathbf{q}_{1},\mathbf{q}_{2},\mathbf{q}_{3}\}$ and $\mathbf{V}=\{\mathbf{v}_{1},\mathbf{v}_{2}\}$ in a high-dimensional spaces, the Hausdorff distance $Haus(\mathbf{Q},\mathbf{V})$ is calculated through the following steps: 1) calculate the maximum of the minimum distances from each vector in $\mathbf{V}$ to all vectors in $\mathbf{Q}$, denoted as $\max_{\mathbf{v}\in\mathbf{V}}\min_{\mathbf{q}\in\mathbf{Q}}d(\mathbf{v},\mathbf{q})=2$; 2) calculate the maximum of the minimum distances from each vector in $\mathbf{Q}$ to all vectors in $\mathbf{V}$, expressed as $\max_{\mathbf{q}\in\mathbf{Q}}\min_{\mathbf{v}\in\mathbf{V}}d(\mathbf{q},\mathbf{v})=3$; 3) determine the Hausdorff distance by $max$ aggregating the results from steps 1 and 2 to obtain $Haus(\mathbf{Q},\mathbf{V})=3$.
Hausdorff distance calculation involves pairwise distance computations between vectors from two sets, followed by aggregation as described in Definition 4. The computational complexity for the Hausdorff distance is $O(m^{2}\cdot d)$, where $m$ is the number of vectors per set and $d$ is the vector dimensionality. This quadratic dependence on $m$ makes exact computation prohibitive for large-scale datasets, especially as set sizes and dimensionality $d$ increase.
@@
To reduce computational overhead, extensive single-vector search works have validated the efficacy of approximation techniques such as quantization [8, 37, 47] and index pruning [24, 15, 28, 14, 20, 10]. Inspired by these, we introduce the approximate top-$k$ vector set search problem.
@@
Approximate top-$k$ vector set search trades off speed and accuracy, permitting a small error in exchange for substantially improved search efficiency.
IV Proposed BioVSS
We propose BioVSS, an efficient search algorithm designed to address the approximate top-$k$ vector set search problem. Drawing inspiration from the fly olfactory circuit, BioVSS leverages the locality-sensitive property of the olfactory circuit to enhance search efficiency. Section IV-B establishes theoretical foundations for locality-sensitive Hausdorff distance.
IV-A Overview Algorithm
BioVSS algorithm comprises two principal components: 1) the binary hash encoding for vector sets, and 2) the search execution in BioVSS.
TABLE I: Summary of Major Notations

### v2

# Approximate Vector Set Search Inspired by Fly Olfactory Neural System

@@ Abstract
Vector set search is an underexplored paradigm of similarity search that focuses on retrieving vector sets similar to a given query set. This paradigm captures the intrinsic structural alignment between sets and real-world entities, enabling finer-grained and more consistent relationship modeling across a wide range of applications. However, it faces significantly greater efficiency challenges than traditional single-vector search, primarily due to the combinatorial explosion of set-to-set comparisons. In this work, we address both the combinatorial complexity inherent in vector set comparisons and the curse of dimensionality carried over from single-vector search. To overcome these obstacles, we propose an efficient algorithm for vector set search, BioVSS  [^1] This work is an extended version of our paper titled Approximate Vector Set Search: A Bio-Inspired Approach for High-Dimensional Spaces, presented at the 41st IEEE International Conference on Data Engineering (ICDE 2025). (Bio-inspired Vector Set Search). BioVSS draws inspiration from the fly olfactory neural system, which employs sparse coding principles for efficient odor recognition in biological neural networks. Based on this biological mechanism, our approach quantizes vectors into sparse binary codes and employs a Bloom filter-based index leveraging set membership properties. Experimental results show that BioVSS achieves over 50× speedup compared to brute-force linear scanning on million-scale datasets, while maintaining a high recall of up to 98.9%, demonstrating its effectiveness for large-scale vector set search.
keywords:
Similarity Search , Vector Set Search , Bloom Filter
journal: x
@@
[whu-cs]organization=School of Computer Science, Wuhan University, city=Wuhan, postcode=430072, state=Hubei, country=China
[…]
[whu-bdi]organization=Big Data Institute, Wuhan University, city=Wuhan, postcode=430072, state=Hubei, country=China
[…]
[amazon]organization=Amazon.com, Inc., city=Seattle, postcode=98109, state=WA, country=USA
1 Introduction
Vector search constitutes a core computational task in numerous fields, including information retrieval [1], recommender systems [2], and computer vision [3]. Most existing methods focus on single-vector queries and aim to address challenges such as the curse of dimensionality [4] and the sparsity of high-dimensional vector spaces [5]. In this work, we investigate a novel and challenging problem termed Vector Set Search. Given a query vector set $\mathbf{Q}$, the goal is to retrieve the top-$k$ most relevant vector sets from a database $\mathbf{D}=\{\mathbf{V}_{1},\mathbf{V}_{2},\dots,\mathbf{V}_{n}\}$. This problem extends beyond conventional single-vector search [6] by involving intra-set vector matching and aggregation, leading to increased computational complexity. At present, there are no efficient solutions available to handle this complexity effectively.
Applications. Vector set search supports a wide range of applications due to its inherent structural alignment with real-world entities, including:
  • Academic Entity Discovery. Academic entity discovery supports efficient navigation of the rapidly expanding scholarly landscape. By modeling academic entities as vector sets—such as representing a paper’s profile [7] through vectors corresponding to its citations, or characterizing an author’s profile [8] as a vector set derived from their publications—vector set search facilitates the identification of related research and the analysis of academic trends.
  • Multi-Modal Query Answering. Multi-modal query answering systems benefit from vector set search to achieve more precise retrieval. As illustrated in [9], multi-modal inputs (e.g., text, images, and audio) are encoded into vector sets using a unified multi-modal encoder [10]. Vector set search then enables cross-modal retrieval by aligning semantic representations across modalities, thereby enhancing the accuracy of the query results.
  • Multi-Vector Ranking for Recommendation. In recommendation systems, multi-vector ranking leverages vector set representations to model users. As shown in [11], users are represented as vector sets that encode different facets of user behavior—such as browsing activities, search histories, and click interactions. This vector set-based ranking framework enables flexible matching across diverse user preferences.
@@ 1 Introduction
Let us delve into scholar search, a characteristic example of vector set search that we examine as our main experimental focus.
Example 1: Figure 1 presents a representative data flow of vector set search in the context of scholar search [8], following the construction of a vector set database. The database preparation stage involves encoding scholar profiles as vector sets based on their publication records [12], forming a structured database of scholars. The vector set search process consists of two primary phases: distance computation and neighbor retrieval. In the distance computation phase, pairwise distances (e.g., Euclidean distance) between the query vector set and each candidate set in the database are computed and subsequently aggregated (e.g., via $max$ and $min$ operations) into a set-level distance. The neighbor retrieval phase then identifies the top-$k$ most similar scholars according to these aggregated distances. This procedure facilitates the discovery of relevant scholars to support downstream research tasks.
Motivation. The preceding example highlights an important observation: the structure of vector sets exhibits strong correspondence with real-world entities, enabling practical and effective applications. With the recent progress in embedding models [12] capable of generating high-dimensional, semantically rich representations, the range of applicable scenarios has significantly broadened. Nonetheless, the computational complexity arising from high-dimensional vectors and the combinatorial growth of pairwise comparisons [13] presents notable challenges for building efficient search solutions.
Recent works such as DESSERT [14] have adopted hash table-based techniques to accelerate certain types of non-metric distance calculations. However, accurately assessing vector set similarity requires more than speed—it demands reliable discrimination and consistency across comparisons. These fundamental properties hinge on the choice of distance metrics (analyzed in Section 3.2). The Hausdorff distance [15, 16], by design, satisfies these criteria, offering a mathematically sound means for evaluating set-to-set similarity while preserving essential symmetric properties critical for dependable comparisons.
Driven by these foundational needs in set-level similarity measurement, we emphasize the importance of using metric distances for maintaining evaluation consistency. The Hausdorff distance, rooted in classical set theory [15, 16], inherently supports set-based distance computation. However, its reliance on exhaustive pairwise distance evaluation renders it computationally expensive, reinforcing the necessity of more efficient approaches for vector set search.
Challenges. Enabling efficient vector set search in high-dimensional settings under Hausdorff distance introduces several core challenges. The first is the well-known curse of dimensionality [4]. While some recent approaches attempt to mitigate this issue through hash table construction [14] or by reducing vector sets to single-vector representations [17], these techniques are inherently limited to specific non-metric distances, thereby restricting their suitability for Hausdorff-based computations. The second challenge lies in the aggregation complexity. Existing algorithms for Hausdorff distance search [18, 19, 20] typically exploit geometric properties to reduce the cost of aggregation. However, such geometric heuristics become ineffective as dimensionality increases, limiting their practical value in high-dimensional scenarios.
Contributions. To overcome the aforementioned challenges, we propose an approximate top-$k$ vector set search algorithm. Our method is inspired by locality-sensitive hashing techniques [21, 22], which are in turn motivated by the biological olfactory circuit [23]. By emulating the mechanisms of the olfactory circuit, we quantize input vectors into sparse binary codes, thereby mitigating the curse of dimensionality. On top of this, we design an indexing structure based on Bloom filters, leveraging their set membership properties to narrow down the search space and reduce the cost of aggregation. Our main contributions are outlined below:
@@
  • We formulate the first approximate vector set search problem in high-dimensional spaces using Hausdorff distance, a natural metric for measuring set similarity (see Section 3).
@@
  • We introduce BioVSS, an algorithm that exploits locality-sensitive properties to accelerate vector set search, and provide formal theoretical analysis with proofs to establish its correctness (see Section 4).
@@
  • We further develop BioVSS++, an improved version of BioVSS that incorporates a dual-layer cascaded filtering mechanism combining inverted index and vector set sketches to minimize unnecessary candidate scans (see Section 5).
@@
  • We perform comprehensive experiments demonstrating that our method achieves more than 50× speedup over linear scan baselines on million-scale datasets, while maintaining a recall rate of up to 98.9%, confirming its practical efficiency (see Section 6).
2 RELATED WORK
@@
Single-Vector Search.  With the advancement of embedding models, single-vector search has emerged as a widely adopted unified paradigm [24]. Nonetheless, achieving high efficiency remains a significant bottleneck in real-world applications. To this end, numerous approximate search strategies [25] have been introduced to accelerate single-vector retrieval. These strategies are broadly classified into three categories: locality-sensitive hashing [22, 26, 27], graph-based techniques [28], and space partitioning schemes [29, 30]. LSH approaches like FlyHash [22] aim to assign similar vectors into identical hash buckets. Graph-based solutions, such as HNSW [28], create navigable proximity graphs to support rapid query traversal. Partitioning-based methods, including IVFFLAT [24, 31], IndexIVFPQ [32, 33, 31], and IVFScalarQuantizer [31], utilize k-means clustering [34] to divide the space and construct inverted indices for faster search. However, these techniques are tailored specifically for single-vector queries, and are not designed to accommodate set-based scenarios.
@@
Vector Set Search.  Vector set search has garnered relatively limited attention and remains underexplored. Earlier studies primarily focused on low-dimensional vector spaces, approaching vector set search as a set matching problem solvable via algorithms such as the Kuhn-Munkres method [35]. In geospatial research, vector sets representing trajectories were treated as point collections and measured using Hausdorff distance [36, 18, 37]. Techniques were proposed to improve efficiency, such as integrating road network structures [38], applying branch-and-bound algorithms [20], and estimating Hausdorff distances for speedup [19]. Nevertheless, these solutions were designed for low-dimensional contexts and do not translate effectively to high-dimensional domains. The rise of embedding models allows vector sets in high-dimensional spaces to express rich semantic information. Some approaches convert a set of vectors into a single vector [17], enabling the application of existing approximate search methods developed for single-vector queries. However, such conversion depends on specific distance functions defined by the corresponding methods. The approach in [14], for instance, constructs a hash table to accelerate vector set search, but has limitations when used with certain metrics. Particularly, it lacks theoretical guarantees for Hausdorff distance when the outer aggregation involves the $max$ function. Moreover, current methods provide insufficient support for Hausdorff distance. Our work addresses this gap by focusing on vector set search in high-dimensional space under the Hausdorff distance metric.
Bio-Inspired Hashing & Bloom Filter.  Biological olfactory systems across species demonstrate exceptional accuracy in odor recognition [39]. For example, mice (rodents) differentiate odors by interpreting neural signals in the olfactory bulb [40], while zebrafish (teleosts) detect amino acids in water through olfactory signaling [41]. Among them, research on flies’ olfactory mechanisms is the most thorough, inspiring the design of sparse locality-sensitive hashing (LSH) functions [21, 22] via computational modeling. Fly-inspired hashing [21], grounded in the fly olfactory system, offers an LSH technique for single-vector search. In this model, specific odor molecules activate only a small subset of neurons, resulting in sparse activations that facilitate robust odor identification. BioHash [22] enhances performance by learning internal data structures to assign connection weights among projection neurons, emulating the biological connectivity of the fly’s olfactory circuit [42]. On a similar note, Bloom filters support efficient set representation by mapping elements to sparse arrays via hash functions. The binary Bloom filter [42] employs a bit array with binary states, while the counting Bloom filter [43] refines this with count-based granularity. This work integrates the fly olfactory circuit’s locality-sensitive behavior with the structural advantages of Bloom filters. We utilize these parallels to build an efficient indexing structure.
3 DEFINITIONS & PRELIMINARIES
This section presents the problem formulation and analyzes why the Hausdorff distance is suitable for vector set comparison.
3.1 Problem Definition
@@
To clarify the computational steps involved in calculating the Hausdorff distance, we present a concrete example below.
Example 2: Figure 2 demonstrates the procedure for computing the Hausdorff distance. Consider two finite vector sets $\mathbf{Q}=\{\mathbf{q}_{1},\mathbf{q}_{2},\mathbf{q}_{3}\}$ and $\mathbf{V}=\{\mathbf{v}_{1},\mathbf{v}_{2}\}$ situated in a high-dimensional space. The Hausdorff distance $Haus(\mathbf{Q},\mathbf{V})$ is obtained via the following operations: (1) compute the maximum over the minimal distances from each vector in $\mathbf{V}$ to all vectors in $\mathbf{Q}$, which yields $\max_{\mathbf{v}\in\mathbf{V}}\min_{\mathbf{q}\in\mathbf{Q}}d(\mathbf{v},\mathbf{q})=2$; (2) compute the maximum over the minimal distances from each vector in $\mathbf{Q}$ to all vectors in $\mathbf{V}$: $\max_{\mathbf{q}\in\mathbf{Q}}\min_{\mathbf{v}\in\mathbf{V}}d(\mathbf{q},\mathbf{v})=3$; (3) aggregate the two values using the $max$ operator to obtain the final Hausdorff distance: $Haus(\mathbf{Q},\mathbf{V})=3$.
The computation of the Hausdorff distance entails evaluating all pairwise distances between vectors in the two sets, followed by aggregation as defined in Definition 4. This leads to a time complexity of $O(m^{2}\cdot d)$, where $m$ denotes the number of vectors in each set and $d$ indicates the dimensionality. The quadratic growth with respect to $m$ renders exact computation impractical for large datasets, particularly when both the set size and dimensionality $d$ are high.
@@
To mitigate this computational burden, prior work in single-vector search has demonstrated the effectiveness of approximation strategies, including quantization [21, 22, 26] and index pruning [29, 30, 28, 32, 33, 31]. Motivated by these insights, we formulate the approximate top-$k$ vector set search problem.
@@
Approximate top-$k$ vector set search achieves higher efficiency by allowing a minor loss in accuracy, effectively balancing computational speed and result precision.
3.2 Suitability Analysis of Hausdorff Distance
Figure 3: Comparative Analysis of Different Distance Measures
The distance measures are defined as follows:
  1. 1.
    Minimum (Min) distance:
  2. 2.
    Mean Minimum (MeanMin) distance:
  3. 3.
    Hausdorff distance: As defined in Definition 4.
[…]
The analysis yields the following results:
  1. 1.
    $d_{min}(\mathbf{Q},\mathbf{A})=d_{min}(\mathbf{Q},\mathbf{B})=1$
  2. 2.
  3. 3.
    $d_{H}(\mathbf{Q},\mathbf{A})=3$, $d_{H}(\mathbf{Q},\mathbf{B})=2$
[…]
The analysis yields the following results:
  1. 1.
    $d_{min}(\mathbf{Q},\mathbf{A})=d_{min}(\mathbf{A},\mathbf{Q})=1$
  2. 2.
  3. 3.
    $d_{H}(\mathbf{Q},\mathbf{A})=d_{H}(\mathbf{A},\mathbf{Q})=3$
4 Proposed BioVSS
We propose BioVSS, a search algorithm optimized for the approximate top-$k$ vector set search task. Inspired by the fly’s olfactory circuitry, BioVSS exploits its inherent locality-sensitive characteristics to improve search performance. The theoretical basis for the locality-sensitive Hausdorff distance is laid out in Section 4.2.
4.1 Overview Algorithm
BioVSS consists of two main components: 1) the binary hash encoding for vector sets, and 2) the search execution in BioVSS.
Table 1: Summary of Major Notations